OpenStudy (anonymous):

Find the solution of 9a2 – 12a – 1 = 0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@DominicNg

OpenStudy (anonymous):

\[9a^2-12a=1\]\[(3a-2)^2=1+4=5\]\[3a-2=\pm \sqrt{5}\]\[a=\frac{ 2\pm \sqrt{5} }{ 3 }\]

OpenStudy (anonymous):

alternatively you can use the quadratic root formula

OpenStudy (anonymous):

thnx